OpenStudy (fandude727):

Identify the number as rational or irrational. Explain. Type your answers below. -√127

OpenStudy (fandude727):

I'm not sure how to do this.

OpenStudy (calculusxy):

Rational numbers can be written as a fraction (like a ratio) Irrational numbers cannot be written as a fraction. Unless the square root of \(-\sqrt{127}\) is a perfect square, then it is irrational.

OpenStudy (fandude727):

So, −√127 would be an irrational number?

OpenStudy (calculusxy):

For example, \(\sqrt{4}\) is a rational number because the square root is equal to 2. \(\sqrt{6}\) is an irrational number thus it needs to be simply left as \(\sqrt{6}\) (unless the direction asks for any approximations).

OpenStudy (calculusxy):

Yes that is correct.

OpenStudy (fandude727):

Alright thank you. you cleared that up for me.

OpenStudy (calculusxy):

No problem :)